# Correlation-induced phase shifts and time delays in resonance enhanced high harmonic generation from Cr+

### v1

@@ Abstract
We investigate resonance-enhanced high harmonic generation (rHHG) in Cr+ by comparing a 1D shape-resonant model, time-dependent density functional theory (TDDFT), and independent particle approximation (IPA) simulations. Previous studies linked rHHG to the $3p\rightarrow 3d$ giant resonance, suggesting a modified four-step model where recolliding electrons are first captured in the autoionizing state before recombining into the ground state, potentially leading to an emission delay associated with the resonance lifetime. While both the 1D-model and TDDFT reproduce experimental spectra, TDDFT reveals that rHHG counterintuitively originates from spin-down $3p$ states, while spin-up $3d$ electrons negligibly contribute. The IPA fails to reproduce rHHG, highlighting the significance of electron correlations. Furthermore, TDDFT revealed a correlation-induced $\sim$480 attosecond time delay, accompanied by a strong phase shift across the resonance, potentially explaining earlier RABBIT measurements. Our work sheds light on long-standing open questions in rHHG and should advance novel ultrafast spectroscopies of electron correlations and resonances.
I Introduction
@@
High-harmonic generation (HHG) is a fundamental process in strong-field physics that enables the generation of coherent extreme ultraviolet (XUV) and soft X-ray radiation[1, 2, 3, 4, 5]. It presents a powerful tool for probing ultrafast electron dynamics and structural properties of matter across phases [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In gases, HHG is commonly well described by a semi-classical three-step model [22, 23] whereby an electron: (i) tunnel ionizes due to intense laser irradiation, (ii) is accelerated by the laser in the continuum, and (iii) recombines with the parent ion, emitting a high-energy photon. The coherence of the electronic wave function throughout this process is key towards utilizing HHG for ultrafast spectroscopy, as has been demonstrated for charge migration[16, 24], molecular symmetry and orientation [25, 26, 27, 28, 29, 30, 31, 32, 33], and more.
The typical HHG spectrum consists of odd harmonics with nearly uniform intensity, forming a characteristic plateau. This uniformity is observed across material systems and arises from the periodicity of the generation mechanism. Consequently, the electronic structure of the generating medium is not directly apparent within the HHG spectrum.
@@
One known gas-phase system that deviates from this three-step model picture is transition-metal ions. Experimentally, HHG from laser-ablated plasma is known to lead to a sharp resonance in which a single harmonic order is significantly enhanced [34, 35, 36]. Resonant HHG (rHHG) has been established to originate microscopically at the single atom level, with macroscopic effects playing a negligible role [36]. Here, the energy of the enhanced harmonic photon aligns with a transition to an existing autoionization state (AIS) in the cation, related to the $3p\rightarrow 3d$ giant resonance (which is a dominant feature in absorption spectra of transition metal ions). This phenomenon involves the excitation of a $3p$ core electron to an unoccupied $3d$ subshell [37, 38, 39]. Its significant strength arises from large electric dipole matrix elements between the $3p$ and $3d$ orbitals, which results from their close spatial overlap. This overlap is strong because the occupied 3d subshell wave function is located inside an effective potential barrier, while higher energy $3d$ orbitals are pushed out due to a centrifugal barrier component of the effective radial potential. Following excitation, the resulting highly excited state (e.g., $3p^{5}3d^{n+1}$) is unstable and rapidly decays through autoionization into continuum states, typically filling the $3p$ hole and emitting an outer electron (e.g., $3p^{6}3d^{n-1}+\epsilon l$). This decay mechanism gives the resonance a large width. The interaction and interference between this indirect autoionization channel and direct photoionization channels (such as $d\rightarrow\epsilon f$) results in the characteristic asymmetric Fano-type line shapes observed in the absorption and photoionization spectra. For Cr+, the ground state configuration is $3p^{6}3d^{5}(^{6}S)$ in terms of total angular momentum notation (LS coupling). The dominant excitation is still the $3p\rightarrow 3d$ transition, i.e. Cr+ $3p^{6}3d^{5}(^{6}S)+\gamma$ Cr+ $3p^{5}(^{2}P)3d^{6}(^{5}D)(^{6}P)$ followed by autoionization into Cr ${}^{2+}\leavevmode\nobreak\ 3p^{6}3d^{4}(^{5}D)+\epsilon l$ with $l=3$ being the dominant channel [37, 39, 38].
In HHG conditions these light-induced transitions are further complexified because electrons can absorb multiple photons simultaneously, allowing multiphoton transitions as well as tunneling to occur. An extended semi-classical 4-step model was suggested to explain rHHG. Essentially, the normal recombination step in the 3-step model is replaced by a capture process in the AIS, followed by a resonant single-photon transition to the ground-state, emitting an enhanced harmonic photon. This can be considered the reverse process of autoionization, in which an energetic photon is first absorbed in the atom and promotes a $3p$ electron to the $3d$ level, followed by the ejection of an electron to the continuum while filling the $3p$ hole.
@@
Numerical simulations for rHHG divide into two categories: (i) model simulations where the AIS is approximated by a shape resonance in a barrier-well potential system, forming a metastable state (MS) [40, 41, 42]. This model correctly reproduces measured enhancements, but does so by fitting the transition energy to experimental observations. With this approach, it remains unclear to what extent the single-electron approximation captures the full dynamics of rHHG, and especially, if there is any time-delay or phase shift incurred to the resonant photon due to passage through the AIS (with reference to RABBIT experiments that suggested a strong phase shift within the resonant harmonic [43]). Approach (ii) is to perform ab-initio simulations of rHHG using either time-dependent density functional theory (TDDFT) [44, 45], or multi-configuration time-dependent approaches [46, 47]. These schemes lead to spectra in agreement with experiments without the need of external parameters, but are more difficult to analyze and extract the physical mechanism. Still, with ab-initio schemes it was established that in some systems correlations play a dominant role in enabling resonant emission [44], and that deeper orbitals can often play a crucial role in dynamics [46]. What remains unclear is how these effects transpire, and how they depend on the electronic spin in magnetic ions such as Cr+. In particular, the phase and time delay of rHHG has not yet been explored or compared with simpler models.
@@
Here we perform an extensive theoretical analysis of rHHG from the magnetic Cr+ system. We employ three complementary levels of theory: (i) a simple shape-resonant model, (ii) ab-initio spin-TDDFT simulations, and (iii) simulations on par with the spin-TDDFT system but where correlations between electrons are artificially frozen, denoted as the independent particle approximation (IPA). This comparison allows us to pin-point the exact mechanisms of rHHG - deeper $3p$ orbitals that generally have a much higher ionization potential and are not expected to contribute. We show that the role of electron correlations is to allow $3p$ electrons to transition into empty $3d$ bands, from which recombination through the AIS is possible. This transition does not occur in uncorrelated theories, and further incurs a time-delay and phase shift onto the rHHG emission, suggesting correlated corrections to the 4-step model. Importantly, we find that this transition is only facilitated by spin-down electrons that can pass through the empty $d$-bands. Our results shed light onto fundamental rHHG physical and could lead to HHG-spectroscopy of attosecond correlated multi-electron systems.
II Results
As a first step, we calculate HHG spectrum and corresponding time-frequency plots from the 1D model under typical experimental conditions. The results in Fig. 1 show a good match to known experimental rHHG spectra from Cr+. We begin by examining the behavior of H29 for different MS lifetimes (Fig. 1 b). Following ref. [48], the lifetime is calculated from the imaginary part of the energy, $\tau=-\frac{\hbar}{2\text{Im}\left[E_{res}\right]}$ We plot the harmonic spectra normalized to the intensity of H27, and use this to quantify the enhancement factor of the resonant H29, defined as $\frac{I_{29}}{I_{27}}$. While the overall harmonic yield decreases with increasing MS lifetime (see SI), the enhancement factor increases. This decrease in yield is consistent with a recollision-based model, since the increase in lifetime is achieved by widening the barrier, which decreases the tunneling rate. To further investigate this interpretation, we repeated the simulation while gradually extending the PML inwards (see the Computational Methods section below and SI) from the box boundary toward its center. As the PML approached the region associated with the semiclassical long trajectories, these trajectories progressively disappeared from the time-frequency plots. In contrast, the short trajectories remained unaffected as they do not extend as far from the nucleus as the long trajectories (see SI). The harmonic spectra was also affected by the advance of the PML into the simulation domain, showing a decrease in yield, beginning with the highest harmonic orders and progressively impacting lower ones. Crucially, the intensity of H29 also decreased. This validates that rHHG in the 1D model is a tunneling-based phenomena, as expected from the 4-step model.
@@ II Results
The time-frequency analysis for an exemplary case is shown in Fig. 1(c). Notably, the electron trajectories calculated using the semi-classical 3-step model (see SI for details) match the simulation with high accuracy. Specifically, there is no distinguishable time delay between the emission time of the resonant H29 and that predicted by the 3-step model. We note that according to the 4-step model, H29 is emitted after the electron is captured in the AIS, therefore, some additional delay is anticipated due to the time the electron spends in the AIS. However, no significant delay is observed in our 1D simulations for varying resonance lifetimes.
@@
Next, we employ the ab-initio TDDFT scheme. Figure 2 shows HHG spectra and time-frequency analysis under similar conditions to Fig. 1 separated to spin-up/down channels (Fig. 2(c,d)), and the total spin-summed response (Fig. 2(b)). While both channels contribute to the resonance emission, the spin-down channel response is about an order of magnitude larger than the spin-up channel, and in fact dominates rHHG. This is counter-intuitive if considering the typical physical mechanism of atomic HHG and the energy level diagram in Fig. 2(b) - the spin-up channel includes initially occupied $3d$ states that have a much smaller ionization potential compared to the $3p$ levels. A-priori, this should make their tunnel ionization and subsequent recombination exponentially more efficient. However, the fact that $3p$ spin-down states dominate the resonant response indicates that they become intimately linked to the AIS through interaction with the laser.
Another surprising feature can be seen in the time-frequency plots in Fig. 2 that is overlaid with the corresponding semi-classical three-step model trajectories. First, we notice that the HHG emission time in the spin-up channel largely follows the semi-classical trajectories with ionization potential matching the $3d$ levels. This is expected since the $3d$ spin-up sub-shell is initially occupied and its ionization energy is much lower than the ionization energy of the $3p$ electrons (which are seen to contribute negligibly at higher energies). The most striking feature is that the HHG emission from the spin-down channel also largely follows the classical trajectories initiated from $3d$ electrons, even though the $3d$ states are not initially occupied by spin-down electrons. This is a direct indication that rHHG is involved and induces a $3p\rightarrow 3d$ transition, while this channel is blocked for the $3p$ spin-up electrons due to the $3d$ spin-up bands being already fully occupied.
@@
To further explore this conjecture, we repeat our simulations but invoking the IPA, i.e. freezing electronic interactions altogether. The results are shown in fig. 3 for both spin channels and the total emission, and reflect a prominent effect - upon freezing correlations the resonance enhancement is completely suppressed. Similar results were recently shown in another system[44]. Here however, the spin-up channel HHG response remains largely similar with/without the IPA, while the spin-down HHG response is substantially suppressed and is the source of the disappearance of rHHG in the absence of correlations. Notably, the emission from the spin-down electrons no longer follows the semi-classical trajectories of the $3d$ electrons, proving that the spin-down $3p\rightarrow 3d$ transition is facilitated by dynamical electron correlations even in the tunneling regime. We further note that while the resonance emission line still exists in the IPA spin-down channel, it is orders of magnitude weaker, and has phase shifted and is now aligned with the maximum of the electric field, similar to the perturbative low-order harmonics. In that respect, even though the atomic system still support a resonant state in the absence of correlations, it does not contribute to rHHG since emission from typical HHG channels is more prominent, and the resonance emission mechanism differs from that of the correlated system.
We further note that the spin-down channel HHG emission in the IPA is very similar in nature to the emission profile obtained from the simple 1D model. Both spectra exhibit a strong dip in the lower harmonics followed by a rise that peaks at resonance. In the 1D model, this has been attributed to the decrease in the ability of the returning electron to penetrate the potential barrier [41]. The similar shape in the IPA spin-down emission, and the lack of resonance in spin-up emission, may be indicative of a shape resonance formed by the time-independent electron-electron interaction[39]. Further, the significantly lower harmonic yield due to IPA spin-down electrons vs spin-up is also consistent with the presence of a shape-resonance, similar to the results in Fig. 1(b).
Overall, we conclude that in the IPA the $3p$ down states do not transition into $3d$ states, and the resonance is emitted by pure multi-photon transitions rather than excitation through the continuum following a separate physical mechanism compared to the correlated theory. In contrast, spin-down $3p$ electrons require electronic correlations in order to resonantly emit HHG in Cr+ in experimental settings, following their transition into $3d$ levels, also incurring a time delay. These results are further corroborated by ab-initio simulations for the single-photon absorption and ionization cross sections (see SI), supporting our conclusions that the down spin channel is the main contributor to absorption at the resonance, which only occurs with correlations.
@@
Finally, we explore the phase of the harmonics at all levels of theory with a RABBIT-like simulation (see details in the SI). From the RABBIT calculation we extract the harmonic phase across orders, as well as the RABBIT sideband (SB) visibility per order. The results presented in Fig. 4(a) show a clear conclusion - a sharp phase jump across the resonance is induced only in the correlated theory. This shift is up to $2.3\pi$ through the resonance, and is followed by a reduction in the expected sideband visibility. A vertical shift of the IPA curve to align with the start or end of the correlated curve (dotted gray lines with dot markers) shows excellent agreement between the models, except near the resonance, highlighting the role of electron correlations in determining harmonic phases in rHHG. Note that we have found that differences in the 1D and IPA simulations can be corrected for with group velocity delay corrections (blue line in Fig. 4(a), see SI for details), which renders both theories analogous. The predicted phase jump in rHHG can potentially explain previous measurements of reduced SB visibility to either side of the resonant harmonic [43], since it suggests a sharp phase jump even within the resonant harmonic itself (the $2.3\pi$ shift occurs abruptly over 4 harmonic orders, translating roughly to $0.5\pi$ jump within the resonant harmonic).
@@
To further test this, Fig. 4(b) presents the actual numerically-observed RABBIT SB visibility and the expected visibility due to two-source interference (taking into account the harmonic yields). The normalized visibility (observed/expected) is shown in Fig. 4(c). For both SB28 and SB30 (closest to the resonance), across all models, the expected visibility is larger than the observed visibility (with the exception of SB30 IPA spin-down channel where they are roughly equal). For both SBs, the normalized correlated visibility is drastically lower than the normalized IPA visibility (SB28 - 15.1% vs 27.1%; SB30 5.6% vs 35%). Looking at the spin-resolved components, we see similar behavior, with the most noticeable case being the spin-down component responsible for rHHG in SB30 (4% vs 103.5%). A similar but opposite difference arises between the spin-up component of SB30 (48.5% vs 4.5%), which we attribute to the extremely low harmonic signal in the IPA spin-up channel. Interestingly, the 1D single-electron model yields a normalized visibility similar to that of the spin-down component in the correlated model. This contrasts with the harmonic phase, which more closely resembles that of the IPA model. While this may seem unexpected, it is important to note that the 1D and IPA models are not equivalent. The 1D model employs a synthetic potential supporting a sharp shape resonance, which may differ from the IPA simulations, where this resonance could be less pronounced. Crucially, only the fully correlated theory accurately reproduces the correct harmonic phase near resonance, highlighting the central role of electron correlation in this regime, also sharply reducing SB visibility in correspondence with measurements[43].
@@
Figure 4: rHHG phase and RABBIT analysis. (a) Harmonic phase comparison with varying levels of theory. Yellow and orange curves denote the fully correlated and IPA simulations, respectively. The blue line denotes the group-delay-corrected 1D simulation (see text). Gray dotted lines with dot markers indicate the vertically shifted IPA results, while the gray dotted line with $\times$ markers corresponds to the unadjusted 1D result. (b) Observed and expected SB visibility (for SB28 and SB30) for various levels of theory and spin channel contributions. (c) Normalized visibility (ratio of observed/expected) for both sidebands and various levels of theory. The correlated theory shows much reduced visibility compared to the expected value, corresponding to the strong phase jump in (a).
III Summary and outlook
@@
In summary, we theoretically explored resonance-enhanced high harmonic generation (rHHG) in Cr+ ions by employing three complementary approaches: (i) a 1D single electron model where the resonance is modeled by a shape-resonance, (ii) ab-initio spin-time-dependent density functional theory (spin-TDDFT), and (iii) spin-TDDFT within the independent particle approximation (IPA) with electronic interactions artificially removed.
@@
The 1D model successfully reproduces experimental rHHG spectra, where we showed that the mechanism for rHHG in this case is tunnel-ionization followed by re-collision (without associated time delays). TDDFT also reproduces experimental spectra, but offers a more comprehensive view into rHHG physics — We uncovered that the primary contribution to enhancement originates from spin-down $3p$ states rather than spin-up $3d$ electrons (that intuitively should dominate due to their lower ionization potential). Strikingly, the more dominant spin-down emission is delayed by $\sim$480 attoseconds compared to the spin-up $3d$ emission. These effects contrast with the 1D model and IPA simulations and point towards a mechanism involving electron interaction and spin selectivity, also aligning with the 4-step picture: A $3d$ spin-up electron tunnel ionizes and accelerates in the continuum. Subsequently, it re-collides with the parent ion, exciting a $3p$ spin-down electron (only possible through correlations), trapping both electrons in the autoionizing state. The autoionization level proceeds to coherently decay to the ground state, emitting a resonantly-enhanced photon. Lastly, we explored RABBIT simulations that investigate the harmonic phases and sideband visibility. These simulations showed a sharp phase jump ($\sim 2.3\pi$) across the resonance that occurred only in the correlated theory. Correspondingly, normalized sideband visibility near the resonance was drastically lower in the correlated theory compared to the IPA (especially for the dominant spin-down component), which might explain earlier measurements[43].

### v2

@@ Abstract
We investigate resonance-enhanced high harmonic generation (rHHG) in Cr+ by comparing a 1D shape-resonant model, time-dependent density functional theory (TDDFT), and independent particle approximation (IPA) simulations. Previous studies linked rHHG to the $3p\rightarrow 3d$ giant resonance, suggesting a modified four-step model where recolliding electrons are first captured in the autoionization state before recombining into the ground state, potentially leading to an emission delay associated with the resonance lifetime. While both the 1D-model and TDDFT reproduce experimental spectra, TDDFT reveals that rHHG counterintuitively originates from spin-down $3p$ states, while spin-up $3d$ electrons negligibly contribute. The IPA fails to reproduce rHHG, highlighting the significance of electron correlations. Furthermore, TDDFT revealed a correlation-induced $\sim$480 attosecond time delay, accompanied by a strong phase shift across the resonance, potentially explaining earlier RABBIT measurements. Our work sheds light on long-standing open questions in rHHG and should advance novel ultrafast spectroscopies of electron correlations and resonances.
@@
High-harmonic generation (HHG) is a fundamental process in strong-field physics, enabling generation of coherent extreme ultraviolet (XUV) and soft X-ray radiation[1, 2, 3, 4, 5]. It presents a powerful tool for probing ultrafast electron dynamics and structural properties of matter across phases [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In gases, HHG is well described by the three-step model [22, 23] whereby an electron: (i) tunnel ionizes, (ii) is accelerated by the laser, and (iii) recombines with the parent ion, emitting a photon. The coherence of the electronic wave function throughout this process is key towards utilizing HHG for ultrafast spectroscopy, as has been demonstrated for charge migration[16, 24], molecular symmetry and orientation [25, 26, 27, 28, 29, 30, 31, 32, 33], and more.
@@
One known gas-phase system that deviates from this three-step model picture is transition-metal ions. Experimentally, HHG from laser-ablated plasma is known to lead to a significant resonant enhancement of a single harmonic order [34, 35, 36]. Resonant HHG (rHHG) has been established to originate at the single atom level, with macroscopic effects playing a negligible role [36]. The energy of the enhanced harmonic photon aligns with a transition to an autoionization state (AIS) in the cation, related to the $3p\rightarrow 3d$ giant resonance. Briefly, this autoionization transition involves the excitation of a $3p$ core electron to an unoccupied $3d$ subshell [37, 38, 39], after which the resulting highly excited state (e.g., $3p^{5}3d^{n+1}$) is unstable and decays via autoionization, typically filling the $3p$ hole and emitting an outer electron (e.g., $3p^{6}3d^{n-1}+\epsilon l$). Resonances in photoionization spectra and resonant photoionization have been the subject of extensive study. A complete topical review is provided in [40]. For Cr+, the ground state configuration is $3p^{6}3d^{5}(^{6}S)$ in LS coupling notation, corresponding to a half-filled $3d$ sub-shell with five spin-up electrons. This configuration breaks symmetry between spin-up/down electrons due to an exchange interaction. The dominant excitation is still the $3p\rightarrow 3d$ transition, i.e. Cr+ $3p^{6}3d^{5}(^{6}S)+\gamma$ → Cr+ $3p^{5}(^{2}P)3d^{6}(^{5}D)(^{6}P)$ followed by autoionization into Cr $^{2+}\penalty 10000\ 3p^{6}3d^{4}(^{5}D)+\epsilon l$ with $l=3$ being the dominant channel [37, 39, 38].
To account for enhancement in rHHG in presence of multi-photon transitions, an extended semi-classical single-active electron (SAE) 4-step model was suggested [41]. It replaces the recombination step of the three-step model with electron capture in the AIS, followed by a transition to the ground state and photon emission.
@@
Alternatively, one can describe rHHG fully numerically with methods falling into two categories: (i) SAE models where the AIS is approximated by a resonance in a barrier-well potential, forming a metastable state (MS) [41, 42, 43, 44]. These models qualitatively reproduces the measured enhancement, but do so by synthetically fitting the transition energies. It remains unclear to what extent the SAE approximation captures the full dynamics of rHHG and if any time-delay or phase shift are inflicted on the resonant photon. Approach (ii) is to perform ab-initio simulations of rHHG using either time-dependent density functional theory (TDDFT) [45, 46], or multi-configuration time-dependent approaches [47, 48]. These schemes lead to spectra in agreement with experiments without the need of external parameters, but are more difficult to analyze and extract the physical mechanism. Still, with ab-initio schemes it was established that in some systems correlations play a dominant role in enabling resonant emission [45], and that deeper orbitals often play a crucial role in dynamics [47]. What remains unclear is how these effects transpire, and how they depend on the electronic spin in magnetic ions such as Cr+. In particular, the phase and time delay of rHHG has not yet been explored with multi-electron theories (only with model semi-analytic approaches[42]).
@@
We perform here an extensive theoretical analysis of rHHG from magnetic Cr+ employing three complementary levels of theory: (i) a shape-resonant SAE model, (ii) ab-initio spin-TDDFT simulations, and (iii) simulations on par with the spin-TDDFT but where electron correlations are artificially frozen, denoted as the independent particle approximation (IPA). This comparison allows pin-pointing the mechanism of rHHG - a low-lying $3d$ electron is tunnel ionized and gains high kinetic energy in the laser field. When it re-collides with the parent ion, it promotes a deep-level $3p$ electron via electron-electron interactions into empty $d$-bands, forming the autoionization state, from which an enhanced rHHG photon is emitted. The $3p$ energy level is situated too deep to typically contribute to HHG if electronic interactions are not invoked. We show that this transition, which does not occur in uncorrelated theories, inflicts a time-delay and phase shift onto the rHHG emission, suggesting correlated corrections to the 4-step model. Importantly, we find that this transition is only facilitated by spin-down electrons that can pass through the empty $d$-bands. Our results shed light onto fundamental rHHG physics and could lead to HHG-spectroscopy of attosecond correlated multi-electron systems.
Let us briefly describe our numerical approach and examined set-up. First, we explore HHG and corresponding time-frequency plots from a 1D SAE model that includes a shape resonance matching withe the experimental resonant transition (all numerical details are delegated to the Appendix). The results in Fig. 1 show a good agreement with experimental rHHG spectra from Cr+ [36, 49]. We examine the behavior of resonant H29 for different resonance lifetimes (Fig. 1 b), especially quantifying its enhancement factor, defined as $\frac{I_{29}}{I_{27}}$. In doing so, we hope to resolve whether a re-collision picture is valid in rHHG in the SAE case. While the overall harmonic yield decreases with increasing lifetime (see SI), the enhancement factor increases. This is consistent with a recollision-based picture, since the increase in lifetime is achieved by widening the barrier, thereby decreasing the tunneling rate. To further investigate this interpretation, we gradually extended the complex absorbing potential (CAP) inwards from the box boundary toward its center. As the CAP approached the semiclassical long trajectories, these progressively disappeared from the time-frequency (Fig. S2), while short trajectories were unaffected. By extending the CAP inward the harmonic yields reduce, including H29, further supporting the recollision mechanism for rHHG.
@@
Time-frequency analysis for an exemplary case is shown in Fig. 1(c). The semiclassical trajectories match the quantum SAE simulations with high accuracy — there is no distinguishable time delay for the resonant H29. We note that according to the 4-step model, H29 is emitted after the electron is captured in the AIS, therefore some delay is anticipated, though we could not resolve it in our conditions.
@@
Next, we employ an ab-initio TDDFT scheme (for details see Appendix). Figure 2 shows HHG spectra and time-frequency analysis under similar conditions to Fig. 1 separated to spin-up/down channels (Fig. 2(c,d)), and the total spin-summed response (Fig. 2(b)). While both channels contribute to the resonance emission, the spin-down channel response is about an order of magnitude larger than the spin-up, dominating rHHG. This is counter-intuitive if considering the typical physical mechanism of atomic HHG and the energy level diagram in Fig. 2(b) - the spin-up channel includes initially occupied $3d$ states that have a much lower ionization potential compared to the $3p$ levels. A-priori, this should make their tunnel ionization and recombination exponentially more efficient. However, the fact that $3p$ spin-down states dominate the resonant response indicates that they become intrinsically linked to the AIS through interaction with the laser. In these calculations, the long trajectories are slightly suppressed compared to the 1D case, as is common in ab-initio theory[50], due to the 3D nature of the simulation.
Another surprising feature can be seen in the time-frequency plots in Fig. 2. First, we notice that the HHG emission time in the spin-up channel largely follows the semi-classical trajectories with ionization potential matching the $3d$ levels. This is expected since the $3d$ spin-up sub-shell is initially occupied and its ionization energy is much lower than the ionization energy of the $3p$ electrons. The most striking feature is that the HHG emission from the spin-down channel also largely follows the classical trajectories initiated from $3d$ electrons, even though the $3d$ states are not initially occupied by spin-down electrons. This is a direct indication that rHHG induces a $3p\rightarrow 3d$ transition, while this channel is blocked for the $3p$ spin-up electrons, as the $3d$ spin-up bands are fully occupied.
@@
To further explore this conjecture, we invoke the IPA, i.e. freezing electronic interactions altogether. The results are shown in fig. 3 for both spin channels and the total emission, and reflect a prominent effect - upon freezing correlations the resonance enhancement is completely suppressed. Similar results were recently shown in another system[45]. Here however, the spin-up channel HHG response remains largely similar with/without the IPA, while the spin-down HHG response is substantially suppressed and is the source of the disappearance of rHHG. Notably, the emission from the spin-down electrons no longer follows the semi-classical trajectories of the $3d$ electrons, proving that the spin-down $3p\rightarrow 3d$ transition is facilitated by dynamical electron correlations even in the tunneling regime. Further, while the resonance emission line still exists in the IPA spin-down channel, it is orders of magnitude weaker, has phase shifted, and is now aligned with the maximum of the electric field, similar to the perturbative low-order harmonics. In that respect, even though the atomic system still supports a resonant state in the absence of correlations, it does not contribute to rHHG since emission from typical HHG channels is more prominent, and the resonance emission mechanism differs from that of the correlated system.
The spin-down emission spectra in both the IPA and correlated models closely resemble the 1D model, exhibiting a dip at lower harmonics followed by a peak at resonance. In the 1D case, this has been attributed to the returning electron’s reduced ability to penetrate the potential barrier [43].
Overall, we conclude that in the IPA, the $3p$ down states do not transition into $3d$ states, and the resonance is emitted by pure multi-photon transitions rather than excitation through the continuum. In contrast, spin-down $3p$ electrons require electronic correlations in order to resonantly emit HHG in Cr+ in experimental settings, following their transition into $3d$ levels, also incurring a time delay. These results are further corroborated by ab-initio simulations for the single-photon absorption and ionization cross sections (see SI), supporting our conclusions that the down spin channel is the main contributor to absorption at the resonance.
@@
Finally, we explore the phase of the harmonics at all levels of theory with a RABBIT-like simulation. We extract the harmonic phase across orders, as well as the RABBIT sideband (SB) visibility per order (see SI for details). The results presented in Fig. 4(a) show a sharp negative phase jump across the resonance that arises only in the correlated theory. This shift is up to $2.3\pi$ through the resonance, accompanied by a reduction in SB visibility. A vertically shifted IPA curve which aligns with the start and end of the correlated curve shows excellent agreement between the models, except near the resonance. We have found that differences in the 1D and IPA simulations can be corrected for with group velocity delay corrections (blue line in Fig. 4(a), see SI for details), rendering both theories analogous. The phase jump in rHHG can potentially explain previous measurements of reduced SB visibility [51], since it suggests a sharp phase jump even within the resonant harmonic itself (the $2.3\pi$ shift occurs abruptly over 3 harmonic orders).
@@
To further test this, Fig. 4(b) presents the actual numerically-observed RABBIT SB visibility and the expected visibility due to two-source interference with unequal intensities. The normalized visibility is shown in Fig. 4(c). For both SB28 and SB30 (adjacent to the resonance), across all models, the expected visibility is larger than the observed with the exception of SB30 IPA spin-down, where they are roughly equal. For both SBs, the normalized correlated visibility is drastically lower than the normalized IPA visibility (SB28 - 15.1% vs 27.1%; SB30 5.6% vs 35%). The spin-resolved components show similar behavior (most noticeably the spin-down component in SB30 (4% vs 103.5%)). An outlier is the spin-up component of SB30 (48.5% vs 4.5%), which we attribute to the extremely low harmonic signal in this channel. Interestingly, the 1D SAE model yields a normalized visibility similar to that of the spin-down component in the correlated model. This contrasts with the harmonic phase, which more closely resembles that of the IPA model. While this may seem unexpected, it is important to note that the 1D and IPA models are not equivalent. The 1D model employs a synthetic potential supporting a sharp shape resonance, which may differ from the IPA simulations, where this resonance could be less pronounced. Crucially, only the fully correlated theory accurately reproduces the correct harmonic phase near resonance, also strongly reducing SB visibility in correspondence with measurements[51].
@@
Figure 4: rHHG phase and RABBIT analysis. (a) Harmonic phase comparison with varying levels of theory. Yellow and orange curves denote the fully correlated and IPA simulations, respectively. The blue line denotes the group-delay-corrected 1D simulation. Gray dotted lines with dot markers indicate the vertically shifted IPA, while the gray dotted line with $\times$ markers corresponds to the unadjusted 1D result. (b) Observed and expected SB visibility (for SB28 and SB30) for various levels of theory and spin channels. (c) Normalized visibility for both sidebands and various levels of theory. The correlated theory shows strongly reduced normalized visibility, corresponding to the strong phase jump in (a).
@@
In summary, we theoretically explored resonance-enhanced high harmonic generation (rHHG) in Cr+ ions by employing three complementary approaches: (i) a 1D SAE model where the resonance is modeled by a shape-resonance, (ii) ab-initio spin-time-dependent density functional theory (spin-TDDFT), and (iii) spin-TDDFT within the independent particle approximation (IPA) with electronic interactions artificially removed.
@@
The 1D model successfully reproduces experimental rHHG spectra, where we showed that the mechanism for rHHG in this case is tunnel-ionization followed by re-collision (without associated time delays). TDDFT also reproduces experimental spectra, but offers a more comprehensive view into rHHG physics — we uncovered that the primary contribution to enhancement originates from spin-down $3p$ states rather than spin-up $3d$ electrons (that intuitively should dominate due to their lower ionization potential). Strikingly, the more dominant spin-down emission is delayed by $\sim$480 attoseconds compared to the spin-up $3d$ emission. These effects contrast with the 1D model and IPA simulations and point to a mechanism involving electron interaction and spin selectivity, also aligning with the 4-step picture: a $3d$ spin-up electron tunnel ionizes and accelerates in the continuum, re-collides with the parent ion, exciting a $3p$ spin-down electron (only possible through correlations), exciting an autoionization state. This proceeds to coherently decay to the ground state, emitting a resonantly-enhanced photon. Lastly, we explored RABBIT simulations that investigate the harmonic phases and sideband visibility. These simulations showed a sharp phase jump ($\sim 2.3\pi$) across the resonance, only in the correlated theory. Correspondingly, normalized sideband visibility near the resonance was drastically lower in the correlated theory compared to the IPA (especially for the dominant spin-down component), which might explain earlier measurements[51].